Let X = {a, b, c} defined by f : X (X such that f = {(a, b), (b, a), (c,c)}.  Find the values of f^1, f^2 and f^4. 


1
Expert's answer
2021-09-20T16:00:29-0400

Solution:

f = {(a, b), (b, a), (c,c)}

"f^{-1} =\{(b,a), ( a,b), (c,c)\}=f"

We found that "f^{-1}=f" for given X.

Then, "\\f^2=f(f)=f(f^{-1})=X"

And "\\f^4=[f(f^{-1})]^2=X^2"